a large rectangular prism measuring 3 cm by 5 cm by 6 cm is painted and then divided up into small cubes measuring 1 cm to a side what is the total unpainted surface area of the small 1 cm^3 cubes?
Answer by ikleyn(52781)

.
a large rectangular prism measuring 3 cm by 5 cm by 6 cm is painted and then divided up
into small cubes measuring 1 cm to a side
what is the total unpainted surface area of the small 1 cm^3 cubes?
~~~~~~~~~~~~~~~~~~

The surface area of the large prism is  2*(3*5 + 3*6 + 5*6) = 126 cm^2.


So, the painted area is 126 cm^2.


The number of small cubes is  3*5*6 = 90.


Their total surface area is 90 times 6 cm^2, i.e. 6*90 = 540 cm^2 (since each small cube has 6 faces).


THEREFORE, unpainted area is 540 - 126 = 414 cm^2.    ANSWER

Solved, answered and explained.